# Circular strings, magnons, plane waves and local quenches in BTZ

### v1

@@ Appendix A Penrose limit of Schwarzschild black hole in $AdS_{5}\times S^{5}$
Next, we go to the Brinkmann Coordinate and calculate the mass matrix
$\tilde{M}_{11}$ $=\left[{E^{2}-J^{2}\left(r^{2}+1-\frac{r_{0}^{2}}{r^{2}}\right)}\right]^{-\frac{1}{2}}\frac{d^{2}}{d\lambda^{2}}\left(\sqrt{1-J^{2}\left(r^{2}+1-\frac{r_{0}^{2}}{r^{2}}\right)}\right)$
$=\frac{d}{dr}\left(\dot{r}\frac{d}{dr}(\dot{r})\right)=-J^{2}\left(1-\frac{3r_{0}^{4}}{r^{4}}\right)$ (286)
$\tilde{M}_{22}$ $=\frac{1}{r}\frac{d^{2}}{d\lambda^{2}}(r)=\frac{1}{r}\left[\dot{r}^{2}\frac{d^{2}}{dr^{2}}+\dot{r}\left(\frac{d\dot{r}}{dr}\right)\frac{d}{dr}\right](r)=\dot{r}\left(\frac{d\dot{r}}{dr}\right)^{2}=-J^{2}\left(1+\frac{r_{0}^{4}}{r^{4}}\right)=\tilde{M}_{33}=\tilde{M}_{44}$ (287)
@@
$ds^{2}=2dx^{+}dx^{-}-J^{2}\left[\left(1-\frac{3r_{0}^{4}}{r^{4}}\right)x_{1}^{2}+\left(1+\frac{r_{0}^{4}}{r^{4}}\right)z_{3}^{2}+z_{4}^{2}\right]\left(dx^{+}\right)^{2}+dz_{3}^{2}+dz_{4}^{2}$ (289)
@@
$F_{5}=R^{4}\left[\text{vol}(AdS_{5})+\text{vol}(S^{5})\right]$ (290)
@@
$\text{Vol}(AdS_{5})=r^{3}dt\wedge dr\wedge\text{Vol}(S^{3}),\quad\text{Vol}(S^{5})=\sin^{4}\psi\ d\psi\wedge\text{Vol}(S^{4})$ (291)
@@
$\text{Vol}(AdS_{5})$ $=R^{4}\ r^{3}dt\wedge dr\wedge\text{Vol}(S^{3})$
$=R^{4}\ r^{3}\left(t_{\lambda}d\lambda-\frac{d\xi}{E}+Jd\phi\right)\wedge\left(r_{\lambda}d\lambda\right)\wedge\text{Vol}(S^{3})$
$=R^{4}\ r^{3}r_{\lambda}\left[-\frac{d\xi}{E}\wedge d\lambda\wedge\text{Vol}(S^{3})+J\ d\phi\wedge d\lambda\wedge\text{Vol}(S^{3})\right]$
$=J\ d\phi\wedge(r_{\lambda}d\lambda)\wedge(r^{3}\text{Vol}(\mathbb{R}^{3}))$ (292)
$\text{Vol}(S^{5})$ $=R^{4}\ \sin^{4}{\psi}\ d\psi\wedge\text{Vol}(S^{4})$
$=R^{4}\ \sin^{4}{(J\lambda+E\phi)}\left\{Jd\lambda+Ed\phi\right\}\wedge\text{Vol}(S^{4})$
$=J\ d\lambda\wedge\sin^{4}{(J\lambda)}\text{Vol}(\mathbb{R}^{4})$
@@
$F_{5}=J\ dx^{+}\wedge\left(dx_{1}\wedge dx_{2}\wedge dx_{3}\wedge dx_{4}+dx_{5}\wedge dx_{6}\wedge dx_{7}\wedge dx_{8}\right)$ (293)

### v2

@@ Appendix A Penrose limit of Schwarzschild black hole in $AdS_{5}\times S^{5}$
Next, we go to the Brinkmann coordinates and calculate the mass matrix
$\tilde{M}_{11}$ $=\left[{E^{2}-J^{2}\left(r^{2}+1-\frac{r_{0}^{2}}{r^{2}}\right)}\right]^{-\frac{1}{2}}\frac{d^{2}}{d\lambda^{2}}\left(\sqrt{1-J^{2}\left(r^{2}+1-\frac{r_{0}^{2}}{r^{2}}\right)}\right),$
$=\frac{d}{dr}\left(\dot{r}\frac{d}{dr}(\dot{r})\right)=-J^{2}\left(1-\frac{3r_{0}^{4}}{r^{4}}\right),$ (286)
$\tilde{M}_{22}$ $=\frac{1}{r}\frac{d^{2}}{d\lambda^{2}}(r)=\frac{1}{r}\left[\dot{r}^{2}\frac{d^{2}}{dr^{2}}+\dot{r}\left(\frac{d\dot{r}}{dr}\right)\frac{d}{dr}\right](r)=\dot{r}\left(\frac{d\dot{r}}{dr}\right)^{2}=-J^{2}\left(1+\frac{r_{0}^{4}}{r^{4}}\right)=\tilde{M}_{33}=\tilde{M}_{44},$ (287)
@@
$ds^{2}=2dx^{+}dx^{-}-J^{2}\left[\left(1-\frac{3r_{0}^{4}}{r^{4}}\right)x_{1}^{2}+\left(1+\frac{r_{0}^{4}}{r^{4}}\right)z_{3}^{2}+z_{4}^{2}\right]\left(dx^{+}\right)^{2}+dz_{3}^{2}+dz_{4}^{2},$ (289)
@@
$F_{5}=R^{4}\left[\text{vol}(AdS_{5})+\text{vol}(S^{5})\right],$ (290)
@@
$\text{Vol}(AdS_{5})=r^{3}dt\wedge dr\wedge\text{Vol}(S^{3}),\quad\text{Vol}(S^{5})=\sin^{4}\psi\ d\psi\wedge\text{Vol}(S^{4}).$ (291)
@@
$\text{Vol}(AdS_{5})$ $=R^{4}\ r^{3}dt\wedge dr\wedge\text{Vol}(S^{3}),$
$=R^{4}\ r^{3}\left(t_{\lambda}d\lambda-\frac{d\xi}{E}+Jd\phi\right)\wedge\left(r_{\lambda}d\lambda\right)\wedge\text{Vol}(S^{3}),$
$=R^{4}\ r^{3}r_{\lambda}\left[-\frac{d\xi}{E}\wedge d\lambda\wedge\text{Vol}(S^{3})+J\ d\phi\wedge d\lambda\wedge\text{Vol}(S^{3})\right],$
$=J\ d\phi\wedge(r_{\lambda}d\lambda)\wedge(r^{3}\text{Vol}(\mathbb{R}^{3})),$ (292)
$\text{Vol}(S^{5})$ $=R^{4}\ \sin^{4}{\psi}\ d\psi\wedge\text{Vol}(S^{4}),$
$=R^{4}\ \sin^{4}{(J\lambda+E\phi)}\left\{Jd\lambda+Ed\phi\right\}\wedge\text{Vol}(S^{4}),$
$=J\ d\lambda\wedge\sin^{4}{(J\lambda)}\text{Vol}(\mathbb{R}^{4}).$
@@
$F_{5}=J\ dx^{+}\wedge\left(dx_{1}\wedge dx_{2}\wedge dx_{3}\wedge dx_{4}+dx_{5}\wedge dx_{6}\wedge dx_{7}\wedge dx_{8}\right).$ (293)